A battery of emf 100v and a resistor of resistance 10 kohm are joined in series. This system is used as a source to supply current to an external resistance R if r is not greater than 100 ohm,the current through it is up to 2 significant digits. What is its value?
1
Expert's answer
2010-07-22T13:11:52-0400
We can find the value of current next way:
I=E/(R+r)=100/(100+10000)= 0 ,0099 A =9.9 mA